Problem 14-10 Prepare a Statement of Cash Flows; Free Cash Flow [LO14-1, L014-2 Joyner Companys income statement for Year 2 follows Sales Cost of goods sold Gross margin Selling and administrative expenses Net operating income Nonoperating items: $ 720,08e 362,888 358,688 218,888 148,988 Gain on sale of equipment Income before taxes Incone taxes Net income 1e,800 15e,8e0 45,800 $ 185,8ee Its balance sheet amounts at the end of Years 1 and 2 are as follows: Year 2 Year 1 Assets Cash Accounts receivable Inventory Prepaid expenses Total current assets Property, plant, and equipnent $ 64,688 $ 89,58e 260,900 11,80 319,gee 277,886 21,988 582,588 1e,see 554,180 633,880e 166,988 514,80 Less accumulated depreciation Net property, plant, and equipnent Loan to Hymans Company Total assets 132,900 66, 188 382e 1,169,208 884,588 49.800 Liabilities and Stockholders Equity Accounts payable Accrued liabilities . $ 317,000 $262,000 57,880 81,5e0 409, 5ee 210 ,0ее 1050e 505,586e 286,008 93,808 15,700 379,888 $1,169, 286 $884,588 42,889 ncome taxes payable Total current 11abilities Bonds payable Total liabilities Conmon stock Retained earnings Total stockholders equity Total liabilities and stockholders equity 443,580 358,ee 165,780

media%2F9f0%2F9f05b851-bca2-44fd-bdc3-bd
Answer #1
STATEMENT OF CASH FLOWS
Cash flows from operating activities:
Net income 105000
Adjustments to reconcile net income to net cash profived by operations:
Depreciation expense (166900-132000) 34900
Gain on sale of equipment -10000
Increase in accounts receivable = 260000-115000 = -145000
Increase in inventory = 319000-277000 = -42000
Decrease in prepaid expenses = 21000-10500 = 10500
Increase in accounts payable = 317000-262000 = 55000
Decrease in accrued liabilities = 57000-42000 = -15000
Increase in income taxes payable = 84500-81500 = 3000 -108600
Net cash provided by operating activities -3600
Cash flows from investing activities:
Sale of equipment 10000
Loan to Hymans Company -49000
Purchase of equipment (633000-514000) -119000
Net cash used in investing activities -158000
Cash flows from financing activities:
Cash from short term note
Issue of bonds (210000-105000) 105000
Dividends (93000+105000-165700) -32300
Issue of common stock (350000-286000) 64000
Net cash from financing activities 136700
Net increase (decrease) in cash -24900
Beginning cash and cash equivalents 89500
Ending cash and cash equivalents 64600
Note:
Details as to original cost and accumulated depreciation on equipment sold is not
available. Hence, the $10000 gain has been shown as sale proceeds in the investment
activities section if the cash flow.